Solve by method of variation of parameters.


d^y/dx^2 + dy/dx + 1 = e^x



1
Expert's answer
2022-03-28T06:19:10-0400

First let’s find a fundamental set of solutions for

"y\u2019\u2019+y\u2019=0"

"y_1=1"; "y_2=e^{-x}"

Then a general solution to the nonhomogeneous differential equation is

"y(x)=-y_1\int\frac{y_2(e^x-1)dx}{W(y_1,y_2)}+""y_2\int\frac{y_1(e^x-1)dx}{W(y_1,y_2)}="

"-\int\frac{(1-e^{-x})dx}{y_1y_2’-y_2y_1’}+""e^{-x}\int\frac{(e^{x}-1)dx}{y_1y_2’-y_2y_1’}=""\int\frac{(1-e^{-x})dx}{e^{-x}}-""e^{-x}\int\frac{(e^{x}-1)dx}{e^{-x}}="

"e^x-x+c_1-e^{-x}(\frac12e^{2x}-e^x-c_2)="

"\frac12e^{x}-x+1+c_1+c_2e^{-x}"

Answer: "y(x)=c_1+c_2e^{-x}+\frac12e^{x}-x+1".